# Kolmogorov–Arnold Networks in Fraud Detection: Bridging the Gap Between Theory and Practice

### v1

@@ 4.2 Genetic Algorithm Results
Table 3: Evaluation Results for Best Parameters Tuned by Genetic Algorithm
@@
4.3 Width Pyramid/Non-pyramid Structure Performance Comparison
The structure of the neural network’s width can significantly affect its performance. A pyramid structure, where each layer has around half the number of nodes as its parent layer, is hypothesized to improve efficiency and accuracy. This section compares the performance of KAN models with pyramid and non-pyramid width structures to determine the most effective configuration.
Results: We ran a full grid search for dataset 1 for hyperparameter tuning. The comparison results are presented, indicating the impact of the width structure on model performance.
@@ 4.3 Width Pyramid/Non-pyramid Structure Performance Comparison
{30, 13, 1} 14.0 4.0 1.00000 0.93636 0.96714
{30, 10, 1} 13.0 5.0 1.00000 0.92727 0.96226
{30, 24, 1} 12.0 4.0 1.00000 0.92727 0.96226
{30, 24, 1} 19.0 4.0 1.00000 0.91818 0.95735
{30, 24, 1} 17.0 4.0 1.00000 0.91818 0.95735
{30, 3, 1} 14.0 30.0 0.91667 0.30000 0.45205
{30, 4, 1} 12.0 3.0 0.97436 0.34545 0.51007
{30, 3, 1} 15.0 30.0 1.00000 0.45455 0.62500
{30, 4, 1} 16.0 3.0 0.96774 0.54545 0.69767
{30, 3, 1} 13.0 30.0 0.55838 1.00000 0.71661
Table 4: Performance Metrics for Different Hyperparameter Configurations of KAN
4.4 Heuristic Search Approach Results
Empirical evidence from our comprehensive grid search for hyperparameters on Dataset 1 facilitated the identification of optimal hyperparameters for the Kernel Attention Network (KAN). By employing a pyramid structure for the width parameter, with $k$ fixed at 15 and the grid parameter set to 5, we aim to achieve relatively good performance across all possible hyperparameters. This methodology not only ensures competitive performance but also significantly reduces computational complexity to $O(1)$.
Results: The best hyperparameter configurations for KAN are presented in Table 5.
@@ 4.4 Heuristic Search Approach Results
Dataset 1 {30, 15, 1} 15 5 1.0 0.89 0.94
Dataset 2 {51, 25, 1} 15 5 0.776 0.757 0.766
Dataset 3 {27, 13, 1} 15 5 0.984 0.692 0.813
Dataset 4 {35, 17, 1} 15 5 0.499 0.879 0.637
Dataset 5 {50, 25, 1} 15 5 0.714 0.624 0.666
Table 5: Performance Metrics of KAN on Datasets 1 to 5
Among all five datasets, only Dataset 1 performed better than most conventional machine learning algorithms. The performance comparison between KAN and other machine learning algorithms for datasets 1 to 5 can be found in the appendix. We also discovered that KAN’s performance is closely related to the dataset’s separability by splines in a 2D space after dimensionality reduction.
@@ 4.5 PCA Results and Quick Decision Rule
Principal Component Analysis (PCA) is a widely used dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional form while retaining most of the variance. This section explores the effectiveness of using PCA to determine the separability of datasets by splines, which can serve as a quick decision rule for the suitability of KAN.
Results: Out of the five datasets, only Dataset 1 is separable by splines with different intervals after being processed by PCA to reduce its dimensionality to 2D, as shown in Figure 2. The results from Kolmogorov–Arnold Networks (KAN) also show that only Dataset 1 performed well. For the other datasets, KAN’s performance was not as good as many conventional machine learning algorithms. This highlights the importance of using PCA as a guidance step to quickly determine whether it is worthwhile to train KAN and the effectiveness of KAN in scenarios where splines can effectively separate the data after dimensionality reduction.
Figure 2: PCA for Dataset 1
Figures 3(a), 3(b), 3(c), and 3(d) illustrate the PCA results for Datasets 2 to 5, respectively. For these datasets, the data was not as effectively separable by splines, which corresponds to the poorer performance of KAN on these datasets.
(a) PCA for Dataset 2
(b) PCA for Dataset 3
(c) PCA for Dataset 4
(d) PCA for Dataset 5
Figure 3: PCA for Datasets 2-5

### v2

@@ 4.2 Genetic Algorithm
Table 3: Best Parameter Configurations Identified by Genetic Algorithm
@@
4.3 Width Structure Comparison
The neural network’s width structure can significantly impact performance. A pyramid structure, where each layer has roughly half the nodes of the previous layer, may enhance efficiency and accuracy. This subsection compares KAN models with pyramid and non-pyramid width structures.
Results: Table 4 shows the impact of width structure on model performance.
@@ 4.3 Width Structure Comparison
{30, 13, 1} 14 4 1.0000 0.9364 0.9671
{30, 10, 1} 13 5 1.0000 0.9273 0.9623
{30, 24, 1} 12 4 1.0000 0.9273 0.9623
{30, 24, 1} 19 4 1.0000 0.9182 0.9574
{30, 24, 1} 17 4 1.0000 0.9182 0.9574
{30, 3, 1} 14 30 0.9167 0.3000 0.4521
{30, 4, 1} 12 3 0.9744 0.3455 0.5101
{30, 3, 1} 15 30 1.0000 0.4545 0.6250
{30, 4, 1} 16 3 0.9677 0.5455 0.6977
{30, 3, 1} 13 30 0.5584 1.0000 0.7166
Table 4: Performance Metrics for KAN Models with Pyramid and Non-Pyramid Width Structures
4.4 Heuristic Search Approach
Our empirical grid search on Dataset 1 allowed us to identify optimal hyperparameters for KAN. By adopting a pyramid structure for the width parameter, fixing $k$ at 15, and setting the grid parameter to 5, we aim to balance performance and computational efficiency, achieving a complexity of $O(1)$.
Results: Table 5 presents the best hyperparameter configurations for KAN across multiple datasets.
@@ 4.4 Heuristic Search Approach
Dataset 1 {30, 15, 1} 15 5 1.0000 0.8900 0.9400
Dataset 2 {51, 25, 1} 15 5 0.7760 0.7570 0.7660
Dataset 3 {27, 13, 1} 15 5 0.9840 0.6920 0.8130
Dataset 4 {35, 17, 1} 15 5 0.4990 0.8790 0.6370
Dataset 5 {50, 25, 1} 15 5 0.7140 0.6240 0.6660
Table 5: Performance Metrics of KAN Across Datasets 1 to 5
Only Dataset 1 outperformed most conventional machine learning algorithms. The comparison between KAN and other algorithms for datasets 1 to 5 is detailed in the appendix. We found that KAN’s performance strongly correlates with the dataset’s separability by splines in a 2D space post dimensionality reduction.
@@ 4.5 PCA Results and Quick Decision Rule
Principal Component Analysis (PCA) is a common dimensionality reduction technique that transforms high-dimensional data into a lower-dimensional form while preserving most of the variance. This subsection explores using PCA to assess the separability of datasets by splines, serving as a quick decision rule for KAN’s suitability.
Results: Among the five datasets, only Dataset 1 was effectively separable by splines after PCA reduced its dimensionality to 2D, as shown in Figure 2. This corresponded with KAN’s superior performance on this dataset. For the remaining datasets, KAN’s performance was not as strong, highlighting the importance of using PCA as a preliminary step in deciding whether to deploy KAN.
Figure 2: PCA Results for Dataset 1
Figures 3(a), 3(b), 3(c), and 3(d) illustrate the PCA results for Datasets 2 through 5. These datasets were not as effectively separable by splines, correlating with KAN’s reduced performance on these datasets.
(a) PCA Results for Dataset 2
(b) PCA Results for Dataset 3
(c) PCA Results for Dataset 4
(d) PCA Results for Dataset 5
Figure 3: PCA Results for Datasets 2-5
5 Conclusion
In this study, we evaluated the potential of Kolmogorov–Arnold Networks (KAN) for fraud detection. While KANs have demonstrated strong performance in specific cases, our research highlights that their utility may not be universally applicable across all fraud detection tasks.
We proposed a practical, quick decision rule using Principal Component Analysis (PCA) to assess whether KAN is a suitable choice for a given dataset. If the data, after dimensionality reduction to two dimensions, can be effectively separated by splines, KAN tends to outperform conventional machine learning models. In scenarios where such separability is not observed, other models might provide comparable or superior performance, making the use of KAN less advantageous.
Furthermore, we introduced a heuristic approach to hyperparameter tuning for KAN, which significantly reduces the computational burden compared to traditional brute-force grid search methods. By adopting a pyramid structure for the width parameter, fixing $k$ at 15, and setting the grid number at 5, we achieved near-optimal performance with a fraction of the computational cost, thereby enhancing the practicality of KAN in real-world applications.
Our findings suggest that while KAN holds promise in certain contexts, its application should be carefully considered based on the nature of the data and the specific requirements of the task. The PCA-based method we proposed offers a valuable tool for quickly assessing KAN’s suitability, ensuring that resources are efficiently allocated to the most appropriate modeling techniques.
Future research could explore the applicability of KAN in other domains and further refine the hyperparameter tuning strategies to enhance its efficiency and effectiveness across a broader range of machine learning tasks.